# mO Haoyu[#1]

### v1

@@ Abstract
Segment Anything Model (SAM) has shown impressive performance in interactive and zero-shot segmentation across diverse domains, suggesting that they have learned a general concept of “objects” from their large-scale training. However, we observed that SAM struggles with certain types of objects, particularly those featuring dense, tree-like structures and low textural contrast from their surroundings. These failure modes are critical for understanding its limitations in real-world use. In order to systematically examine this issue, we propose metrics to quantify two key object characteristics: tree-likeness and textural separability. Through extensive controlled synthetic experiments and testing on real datasets, we demonstrate that SAM’s performance is noticeably correlated with these factors. We link these behaviors under the concept of “textural confusion”, where SAM misinterprets local structure as global texture, leading to over-segmentation, or struggles to differentiate objects from similarly textured backgrounds. These findings offer the first quantitative framework to model SAM’s challenges, providing valuable insights into its limitations and guiding future improvements for vision foundation models.
Introduction
Segment Anything Model (SAM) [27], the most widely-used foundation model for image segmentation, has demonstrated promising zero-shot and fine-tuned segmentation ability for objects not seen in training, within a wide range of domains since its release, such as biomedical images [30, 22, 10], remote sensing [6, 35], etc. However, certain failure modes for SAM have been found empirically, where SAM produces surprisingly underwhelming performance on datasets with specific uncommon objects or atypical contexts, such as those with dense, branching structures (e.g., retinal blood vessels) [36, 11], concealed/low-contrast objects [7, 39, 24], small or irregular objects, and others [24] [^1] We focus on “unintentional” failure modes of SAM, as opposed to intentional adversarial attacks/examples [45, 47].. A solution to this problem could simply be fine-tuning the model on such challenging examples using various strategies [29, 32, 17], but this would still lack an understanding of the underlying causes of the issues, and so may not result in the best solution. As such, it is important to develop a better theory for why these failures happen, in order to better understand (and potentially mitigate) the limitations of SAM for new applications.
@@ Introduction
In this paper, we propose that several of these failure modes can be explained as SAM misinterpreting textural cues for how it disambiguates an object from its surroundings: it typically confuses objects with irregular, dense tree-like local structure as large-scale textures (resulting in over-segmentation), and similarly is generally challenged by objects that have similar textures to their surroundings. We propose metrics developed from first principles for quantifying the characteristics of (1) object tree-likeness and (2) textual separability. Objects tree-likeness (Sec. 2.2) is based on concepts such as spatial contour frequency and the difference in typical object variability between global and local scales. Textural separability (Sec. 3.1) is based on the detectability of differences in textural features of an object compared to its surroundings. We designed these metrics to be simple and interpretable, such that they can be applied to any dataset or object type. We also present them in quickly-computable PyTorch-like algorithmic form that can be applied to any masks.
Next, we begin to probe the effect that these characteristics have on SAM’s segmentation ability through experiments on synthetic data which are designed to cover a range of object tree-likeness (Sec. 2.3.1) and textural separability (Sec. 3.2.1). These experiments are carefully controlled to disentangle the effects of object shape and texture on performance, as well as mitigate any other potential confounding factors. We then continue with similar experiments on real datasets (Secs. 2.3.2 and 3.2.2) that represent a wide range of objects to segment. In all experiments, we find both the exact measured tree-likeness and the textural separability of objects are noticeably correlated with SAM’s performance (IoU) in segmenting it, over a wide range of object and image types. Our findings are the first to quantitatively model and experimentally verify how the performance of segmentation foundation models is affected by certain measurable object characteristics, providing an explanation for prior findings of SAM’s unexpected failures on new datasets.
Our overall contributions are as follows:
@@
    Based on first principles, we propose quantitative metrics for object tree-likeness of Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD), as well as a metric for object textural separability.
@@
    We present carefully-designed experiments for the segmentation of synthetic images which cover a range of tree-likeness and textural separability, where important factors such as object shape and texture are controlled.
@@
    We show that SAM’s segmentation performance tends to negatively correlate with an object’s tree-likeness, and positively correlate with it’s textural separability, on a variety of experiments on both synthetic and real images.
We hope that this inspires future work in understanding the limitations of vision foundation models. We release easy-to-use code for our proposed metrics at https://github.com/mazurowski-lab/SAM-TexturalConfusion-Metrics.
Related Works
@@
It has also been found that the shape-learning behavior of neural networks can differ noticeably depending on the particular training dataset and objective [21, 20], motivating us to study the specific behavior of SAM (given its’ unique training dataset and pipeline) in detail. Interestingly, there has been some evidence that vision transformers (like SAM’s image encoder) may instead be biased towards shape, rather than texture for classification tasks [40]. Here, we study segmentation rather than classification, where it is unclear if those results extend, as our results generally point to SAM focusing on texture over shape (to the point of even confusing complex, locally-varying shape itself as texture).
1 Methods: SAM Usage and Prompting
We experiment with both the ViT-H (default) and ViT-B standard pretrained SAM models [27]. As is default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$. All images that possess instance segmentations for multiple objects (e.g., iShape and Plittersdorf) will be evaluated by SAM segmenting each of the images’ objects one at a time (guided by clear prompting), with all other mask pixels set as background. In all experiments, we use the oracle-chosen mask out of SAM’s three outputted predictions.
For all experiments on the synthetic images of Sec. 2.3.1, we provide a tight bounding box about the object of interest as the prompt. For all experiments on real images as well as the style-transferred images of Sec. 3.2.1, we provide the same bounding box as well as a certain number of positive and/or negative point prompts randomly sampled from the object foreground and background within the bounding box, respectively, depending on the dataset (full details in Appendix C.1). We use this relatively heavy prompting strategy in order to minimize the ambiguity of instructions provided to SAM, which will also help to minimize any differences of the oracle prediction from the other two.
@@ 2.1 Motivation
We first wished to understand SAM’s challenges in segmenting certain objects with dense, tree-like structures, following our observation of a trend of SAM having difficulty with such objects including retinal blood vessels [34, 29] or satellite road images [14, 42] (see e.g., Fig. 2). Interestingly, while both of these object types have branching features, a characteristic which naively could relate to SAM’s failure, SAM’s performance on retinal vessel images is noticeably worse (avg. $\mathrm{IoU}\simeq 0.05$ [34]) compared to on satellite road structures (avg. $\mathrm{IoU}\simeq 0.2$ [14]), which was reproduced in our own experiments (Fig. 4).
@@
This motivated us to quantify the features of these types of objects that relate to SAM’s failure. We hypothesize that such objects become more challenging to segment when their branching structures are dense and irregularly-spaced, which we refer to as “tree-like”. We will show that SAM considers such structures as textures rather than shapes, resulting in significant over-segmentation. In the following section we propose how to characterize the tree-likeness of objects using two new quantitative metrics: CPR (Contour Pixel Rate) and DoGD (Difference of Gini Impurity Deviation).
@@ 2.2.2 Difference of Gini-impurity Deviation (DoGD)
We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values which resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using ViT-H SAM. We show results using a wide range of other values for these hyperparameters in Appendix B.1, where we found our findings to be consistent for most other settings. Moreover, we note that CPR and DoGD are correlated with each other (Pearson $|r|=0.83$ on average; Appendix D.1), showing their consistency in how they quantify different aspects of tree-likeness.
@@ 2.3.1 Experiments on Synthetic Data
In this section, we will first carefully probe the effect of object tree-likeness on SAM’s segmentation performance by testing it on synthetic images which solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures. These objects are contiguous components samples from retinal blood vessel and satellite road masks, with the full procedure of generating these images detailed in Appendix A.2. Example generated images, masks and SAM segmentation predictions for them are shown in Fig. 3. As shown (as well in Fig. 4), the objects cover a wide range of tree-likeness as measured by these quantities.
In order to mitigate any confounding on SAM performance due to an object’s textural contrast from its surroundings (which we study in Sec. 3), for each mask $m_{c}$ generated by our procedure, we apply SAM to $K=7$ images created by applying $K$ different randomly-sampled pairs of textures to the object’s foreground ($m_{c}=1$) and background ($m_{c}=0$). We then obtain SAM’s final prediction for this object via pixel-wise majority voting over its predictions on these $K$ images.
Figure 3: Left: Example synthetic tree-like images and object masks. Right: Trend of increasing tree-likeness (increasing CPR/decreasing DoGD) of these objects resulting in worse SAM segmentation predictions (to the right of each object mask).
In Fig. 4 we show the relationship between an object’s tree-likeness (as measured by CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects. We quantify the strength of this relation via rank/non-linear correlation, measured by Kendall’s tau ($\tau$) [25] and Spearman’s rho ($\rho$) [37], shown in Table 1333We evaluate $\tau$ in addition to $\rho$ due to it being more robust to outliers.. Intriguingly, we see that object tree-likeness is quite predictive of SAM’s performance, with average absolute correlations of $|\tau|=0.76$ and $|\rho|=0.93$ for CPR and $|\tau|=0.64$ and $|\rho|=0.82$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance.
@@
SAM Encoder $\tau$ $\rho$ $\tau$ $\rho$
ViT-H -0.77 -0.93 0.61 0.80
ViT-B -0.75 -0.93 0.66 0.84
Table 1: Rank correlation between SAM segmentation IoU and object tree-likeness (CPR and DoGD), on the synthetic dataset.
Figure 4: SAM segmentation IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset.
@@ 2.3.2 Experiments on Real Data
We will now perform the same analysis on two real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k and iShape. DIS5k [33] (or “DIS” for short) is a dataset that contains extremely detailed segmentation masks of objects with varying degrees of hollowness, and both regular and irregular tree-like structures. All DIS experiments will be reported on its validation set unless otherwise stated. We show example images with objects with varying degrees of tree-likeness (by CPR) in Fig. 1 left.
@@
Results.
In Fig. 5 we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, with accompanying quantitative correlation results shown in Table 2. In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs. textural separability experiments of Sec. 3.
@@ Results.
CPR: ViT-H -0.59 -0.76 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
ViT-B -0.63 -0.81 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
DoGD: ViT-H 0.58 0.77 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
ViT-B 0.45 0.64 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
Table 2: Rank correlation between SAM prediction IoU and object tree-likeness (CPR, upper table) and DoGD, lower table), on DIS and iShape.
Figure 5: SAM segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) and iShape (right).
We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SAM segmentation performance. In particular, we see an average absolute correlations of $|\tau|=0.56$ and $|\rho|=0.74$ for CPR and $|\tau|=0.44$ and $|\rho|=0.62$ for DoGD, excluding the antenna and wire objects of iShape which had outlying correlations, likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig. 5), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.

### v2

# Quantifying the Limits of Segmentation Foundation Models: Modeling Challenges in Segmenting Tree-Like and Low-Contrast Objects

@@ Abstract
Image segmentation foundation models (SFMs) like Segment Anything Model (SAM) have achieved impressive zero-shot and interactive segmentation across diverse domains. However, they struggle to segment objects with certain structures, particularly those with dense, tree-like morphology and low textural contrast from their surroundings. These failure modes are crucial for understanding the limitations of SFMs in real-world applications. To systematically study this issue, we introduce interpretable metrics quantifying object tree-likeness and textural separability. On carefully controlled synthetic experiments and real-world datasets, we show that SFM performance (e.g., SAM, SAM 2, HQ-SAM) noticeably correlates with these factors. We link these failures to “textural confusion”, where models misinterpret local structure as global texture, causing over-segmentation or difficulty distinguishing objects from similar backgrounds. Notably, targeted fine-tuning fails to resolve this issue, indicating a fundamental limitation. Our study provides the first quantitative framework for modeling the behavior of SFMs on challenging structures, offering interpretable insights into their segmentation capabilities.
1 Introduction
Segment Anything Model (SAM) [26], the most widely-used segmentation foundation model (SFM), has demonstrated promising zero-shot and fine-tuned segmentation ability across diverse domains, including biomedical imaging [31, 21, 10], remote sensing [6, 37], and others. However, certain failure modes for SAM have been found empirically, where SAM produces underwhelming performance on datasets with certain uncommon objects or atypical contexts, such as dense, branching structures (e.g., retinal blood vessels) [38, 11], concealed/low-contrast objects [7, 41, 23, 44], small or irregular objects, and others [23].
We will show that newer SFMs, including SAM 2 [36] and HQ-SAM [24], are also susceptible to these failure modes, despite some stated improvements over these limitations for SAM—HQ-SAM on thin, branching structures and SAM 2 on low-contrast objects. Fine-tuning or adapting these models for such challenging examples using various strategies [30, 33, 16, 24] (or utilizing specialized techniques, e.g., matting algorithms [28]) could mitigate these issues, yet we will show that these strategies do not fully resolve the problems, nor explain the root causes. This underscores the need for a theoretical understanding of why SFMs struggle with such cases, paving the way for better model design and adaptation for new applications.
@@ 1 Introduction
Here, we propose that several of these failure modes in SFMs stem from misinterpreting textural cues when distinguishing objects from their surroundings. Specifically, SFMs often confuse dense, tree-like structures as large-scale textures, leading to over-segmentation, and struggle with objects that have similar textures to their surroundings. To quantify these challenges, we introduce interpretable metrics for object tree-likeness and textural separability. Tree-likeness (Sec. 4.2) captures structural complexity using spatial contour frequency and the difference in object variability between global and local scales. Textural separability (Sec. 5.1) measures how distinguishable an object’s textural features are from those of its background. These metrics are designed for broad applicability across datasets and object types and are implemented in a PyTorch-compatible form for efficiency.
We then investigate how tree-likeness and textural separability affect SFM performance through controlled synthetic experiments (Secs. 4.3.1, 5.2.1) and real-world datasets (Secs. 4.3.2, 5.2.2). These experiments systematically isolate the influence of object shape and texture while minimizing confounding factors. Across all datasets, we find that SFM performance (IoU) noticeably correlates with both tree-likeness and textural separability across diverse object and image types. Notably, HQ-SAM, designed for intricate structures, still exhibits these limitations, highlighting the persistence of these challenges. Furthermore, our fine-tuning experiments (Sec. 6) confirm that these failure modes remain robust to fine-tuning, even when explicitly trained on the most difficult cases. Our overall contributions are as follows:
@@
    We introduce quantitative metrics for object tree-likeness—Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD)—and an object textural separability metric based on neural feature contrast.
@@
    Through carefully controlled synthetic experiments and real datasets, we demonstrate that the segmentation performance of SAM, SAM 2, and HQ-SAM negatively correlates with tree-likeness and positively correlates with textural separability.
@@
    We show that these segmentation challenges persist even after fine-tuning, suggesting that SFMs struggle with these object types regardless of adaptation.
Our findings provide the first quantitative model linking SFM performance to measurable object characteristics. Despite the complexity of SFMs and their evaluation datasets, we show that their underperformance on challenging cases can be modeled using our straightforward, interpretable object descriptors. We release easy-to-use code for our proposed metrics at https://github.com/mazurowski-lab/SAM-TexturalConfusion-Metrics.
2 Related Work
@@
It has also been found that the shape-learning behavior of neural networks can differ noticeably depending on the particular training dataset and objective [20, 19], motivating us to study the specific behavior of SAM (given its unique training set) and other similar SFMs in detail. Interestingly, there has been some evidence that vision transformers (like SAM’s image encoder) may instead be biased towards shape, rather than texture for classification tasks [42]. Here, we study segmentation rather than classification, where it is unclear if those results extend, as our results generally point to SAM focusing on texture over shape (to the point of even confusing complex, locally varying shape itself as texture), which aligns with the preliminary findings of [48].
3 Methods: SFM Usage and Prompting
We experiment with the ViT-H (default, if unstated) and ViT-B standard pre-trained SAM models [26], the ViT-L and ViT-B+ variants of SAM 2 [36], and the ViT-H and ViT-B variants of HQ-SAM [24], a SAM adaptation designed particularly for thin structures. As is the default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$. All images that possess instance segmentations for multiple objects will be evaluated by the given SFM segmenting each of the images’ objects one at a time (guided by clear prompting described in the next paragraph), with all other mask pixels set as background. In all experiments, we use the oracle-chosen mask out of the given SFM’s three outputted predictions.
For all experiments on the synthetic images of Sec. 4.3.1, we provide a tight bounding box about the object of interest as the prompt. For all experiments on real images as well as the style-transferred images of Sec. 5.2.1, we provide the same bounding box as well as a certain number of positive and/or negative point prompts randomly sampled from the object foreground and background within the bounding box, respectively, depending on the dataset (full details in Appendix C.1). We use this relatively heavy prompting strategy to minimize the ambiguity of instructions provided to the given SFM, which will also help to minimize any differences in the oracle prediction from the other two.
@@ 4.1 Motivation
We first wished to understand the challenges for SFMs in segmenting certain objects with dense, tree-like structures, and then quantify the features driving these challenges, following our observation of a trend of SAM having difficulty with such objects, including retinal blood vessels [35, 30] or satellite road images [13, 45] (see e.g., Fig. 2). Interestingly, while both of these object types have branching features, a characteristic which naively could relate to the SFM’s “failure”, SAM’s performance on retinal vessel images is noticeably worse (avg. $\mathrm{IoU}\simeq 0.05$ [35]) compared to on satellite road structures (avg. $\mathrm{IoU}\simeq 0.2$ [13]), which was reproduced in our own experiments (Fig. 4) [^1] This could potentially simply be due to retinal vessels being thinner than road structures, but our experiments do not support this (App. D.2)..
@@
We hypothesize that such objects become more challenging to segment when their branching structures are dense and irregularly-spaced, which we refer to as “tree-like”. We will show that SFMs consider such structures as textures rather than shapes, resulting in significant over-segmentation. In the following section, we propose how to characterize the tree-likeness of objects using two new quantitative metrics: CPR (Contour Pixel Rate) and DoGD (Difference of Gini Impurity Deviation).
@@ 4.2.2 Difference of Gini-impurity Deviation (DoGD)
We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values that resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using ViT-H SAM. We show results using a wide range of other values for these hyperparameters in Appendix B.1, where our findings were consistent for most other settings. Moreover, we note that CPR and DoGD are correlated with each other (Pearson $|r|=0.83$ on average; Appendix D.1), showing their consistency in quantifying different aspects of tree-likeness.
@@ 4.3.1 Experiments on Synthetic Data
In this section, we will first probe the effect of object tree-likeness on the performance of SFMs by testing them on synthetic images that solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures. These objects are contiguous components sampled from retinal blood vessel and satellite road masks; the full procedure of generating these images is detailed in Appendix A.2. Example generated images, masks, and SAM predictions for them are shown in Fig. 3. As shown (also in Fig. 4), the objects cover a wide range of tree-likeness as measured by these quantities.
In order to mitigate any confounding effects on a given SFM’s performance due to an object’s textural contrast from its surroundings (which we study in Sec. 5), for each mask $m_{c}$ generated by our procedure, we apply the SFM to $K=7$ images created by applying $K$ different randomly-sampled pairs of textures to the object’s foreground ($m_{c}=1$) and background ($m_{c}=0$). We then obtain the SFM’s final prediction for this object via pixel-wise majority voting over its predictions on these $K$ images.
Figure 3: Left: Example synthetic tree-like images and object masks. Right: Trend of increasing tree-likeness (increasing CPR/decreasing DoGD) of these objects (left in each pair) resulting in worse SAM segmentation predictions (right in each pair)
In Fig. 4 we show the relationship between an object’s tree-likeness (via CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects. We quantify the strength of this relation via non-linear correlation, measured by Kendall’s tau ($\tau$) [25] and Spearman’s rho ($\rho$) [39], for all three SFMs, shown in Table 1. We see that object tree-likeness is quite predictive of an SFM’s performance, with average absolute correlations of $|\tau|=0.75$ and $|\rho|=0.92$ for CPR and $|\tau|=0.67$ and $|\rho|=0.84$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance.
@@
Model $\tau$ $\rho$ $\tau$ $\rho$
SAM, ViT-H [26] -0.77 -0.93 0.61 0.80
SAM, ViT-B [26] -0.75 -0.93 0.66 0.84
SAM 2, ViT-L [36] -0.73 -0.91 0.67 0.85
SAM 2, ViT-B+ [36] -0.71 -0.90 0.69 0.86
HQ-SAM, ViT-H [24] -0.74 -0.92 0.70 0.87
HQ-SAM, ViT-B [24] -0.81 -0.95 0.66 0.84
Average -0.75 -0.92 0.67 0.84
Table 1: Rank correlation between SFM IoU and object tree-likeness (CPR and DoGD), on the synthetic dataset.
Figure 4: SAM IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset.
@@ 4.3.2 Experiments on Real Data
We will now perform the same analysis on two real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k and iShape. DIS5k [34] (or “DIS”) is a dataset containing extremely detailed annotations of objects with varying degrees of hollowness, and both regular and irregular tree-like structures. All DIS experiments will be reported on its validation set unless otherwise stated. We show example images with objects with varying degrees of tree-likeness (by CPR) in Fig. 1 left.
@@
In Fig. 5 we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, alongside the other SFMs (SAM 2 and HQ-SAM) for comparison for DIS, with accompanying quantitative correlation results for all three SFMs shown in Table 2. In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs. textural separability experiments of Sec. 5.
@@
Model $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$
CPR: SAM, ViT-H [26] -0.59 -0.76 -0.13 -0.21 -0.61 -0.81 -0.61 -0.81 -0.25 -0.37 -0.60 -0.77 -0.06 -0.10
SAM, ViT-B [26] -0.63 -0.81 -0.18 -0.27 -0.63 -0.82 -0.63 -0.83 -0.44 -0.62 -0.62 -0.80 -0.12 -0.17
SAM 2, ViT-L [36] -0.41 -0.57 -0.30 -0.43 -0.64 -0.84 -0.61 -0.81 -0.48 -0.66 -0.67 -0.85 -0.14 -0.21
SAM 2, ViT-B+ [36] -0.38 -0.52 -0.16 -0.23 -0.52 -0.71 -0.57 -0.78 -0.42 -0.59 -0.60 -0.77 -0.12 -0.18
HQ-SAM, ViT-H [24] -0.52 -0.69 -0.23 -0.33 -0.77 -0.93 -0.79 -0.95 -0.51 -0.70 -0.69 -0.85 -0.22 -0.33
HQ-SAM, ViT-B [24] -0.62 -0.81 -0.25 -0.34 -0.75 -0.92 -0.83 -0.96 -0.50 -0.69 -0.71 -0.87 -0.24 -0.35
Average -0.53 -0.69 -0.21 -0.30 -0.65 -0.84 -0.67 -0.86 -0.43 -0.61 -0.65 -0.82 -0.15 -0.22
DoGD: SAM, ViT-H [26] 0.58 0.77 0.23 0.30 0.52 0.70 0.46 0.65 0.52 0.72 0.30 0.46 0.08 0.11
SAM, ViT-B [26] 0.45 0.64 0.12 0.14 0.51 0.69 0.45 0.64 0.50 0.70 0.13 0.22 0.06 0.08
SAM 2, ViT-L [36] 0.44 0.60 0.42 0.59 0.48 0.65 0.49 0.68 -0.05 -0.08 0.56 0.75 -0.03 -0.05
SAM 2, ViT-B+ [36] 0.26 0.39 0.40 0.56 0.29 0.40 0.30 0.44 -0.06 -0.10 0.42 0.58 -0.08 -0.12
HQ-SAM, ViT-H [24] 0.48 0.66 0.53 0.72 0.58 0.72 0.48 0.67 -0.31 -0.45 0.57 0.75 -0.13 -0.19
HQ-SAM, ViT-B [24] 0.50 0.69 0.67 0.85 0.59 0.73 0.57 0.77 -0.25 -0.34 0.53 0.71 -0.04 -0.06
Average 0.45 0.63 0.40 0.53 0.50 0.65 0.46 0.64 0.06 0.08 0.42 0.58 -0.02 -0.04
Table 2: Rank correlation between SFM IoU and object tree-likeness (CPR, upper table and DoGD, lower table), on DIS and iShape.
Figure 5: Segmentation IoU vs. object tree-likeness (CPR, top; DoGD, bottom) on DIS (left) for all three SFMs and iShape (right) shown for SAM.
We see that despite the many potential confounding factors in real data, there is still a clear correlation between object tree-likeness as measured by the proposed metrics and SFM segmentation performance. In particular, we see average absolute correlations of $|\tau|=0.59$ and $|\rho|=0.76$ for CPR and $|\tau|=0.38$ and $|\rho|=0.51$ for DoGD—excluding the antenna and wire objects of iShape which had outlying correlations likely because those two object classes cover only small ranges of tree-likeness as opposed to the other iShape classes (Fig. 5), such that noise from other confounding factors obscures any dependence of performance on tree-likeness.